# Half canonical rings of Gorenstein spin curves of genus two

### v1

@@ 1. Introduction
Recall that a spin structure (or theta-characteristic) on a smooth curve $C$ is a line bundle ${\mathcal{L}}$ such that there is an isomorphism $\mu\colon{\mathcal{L}}\otimes{\mathcal{L}}\to\omega_{C}$. The interest in such structures originated in the classical study of theta functions in the 19th century and they were taken up again in 1970’s in the context of the study of moduli spaces of curves (see [Far12]). They also occur in the study of algebraic surfaces: if $X$ is a smooth projective surface and $C\in|K_{X}|$ is a smooth curve in the canonical linear system, then ${\mathcal{L}}=K_{X}|_{C}$ is a theta-characteristic by the adjunction formula.
Both in applications to moduli and to algebraic surfaces it soon became necessary to consider degenerations of the above description, with both the curve and the line bundle possibly acquiring singularities. In the context of moduli of curves this problem has seen some attention [Cor89, Jar98, Lud10]. However, the application to algebraic surfaces, that sparked our interest in these structures and will be explained in a subsequent paper, requires a different approach than what was available in the literature.
More precisely, the curves $C$ we are led to consider are Gorenstein curves $C$ of genus two with ample canonical bundle $\omega_{C}$. Such curves, if reduced, can be classified into two types, Type A and Type B, according to the image of the bicanonical map (Proposition 3.3).
A generalised Gorenstein spin (ggs) structure on $C$ is given by a torsion-free sheaf ${\mathcal{L}}$ everywhere of rank one such that $\chi({\mathcal{L}})=0$ and a multiplication map $\mu\colon{\mathcal{L}}\operatorname{\otimes}{\mathcal{L}}\to\omega_{C}$ which is an isomorphism at the generic points. If ${\mathcal{L}}$ is in fact a line bundle, what we call a regular ggs structure, then this is the direct generalisation of the original notion. We will see in Section 4 that otherwise the ggs structures on $C$ are governed by regular ggs structures on partial normalisations of $C$ in many cases. This leads to a pretty complete understanding of the possible ggs structures on the reduced curves of interest. The non-reduced cases, which presents more complications, is not addressed here.
@@
For stable curves our definition of spin structure coincides with the definition given in [Jar98] but, in view of the application to stable surfaces, we need to include both curves with cusps and curves with elliptic tails in our analysis. As explained in the comments after Definition 1.2 in [FS13], these two classes of curves cannot occur at the same time in a separated moduli problem for curves, so it is not clear how to deal with families in our setup. We have chosen not to pursue this question here.
@@
We then go on and analyse the half-canonical ring of such a ggs structure, that is,
@@
where $R_{2n}=H^{0}(C,nK_{C})$ and $R_{2n+1}=H^{0}(C,{\mathcal{L}}(nK_{C}))$. Note that this is not a standard section ring, because the multiplication of two elements of odd degree has to be defined via the multiplication map $\mu$. Leaving the precise statements to Section 5 we prove the following.
Theorem 5.2 (simplified version) —
Let $({\mathcal{L}},\mu)$ be a ggs structure on a reduced Gorenstein curve of genus two with ample canonical bundle. Then the following cases occur:
@@
In cases $A(0)$, $A(1)$, and $B(2)$ we get complete intersections, in cases $B(0)$ and $B(1)$ the rings admit a more complicated presentation in codimension six and four.
Note that in each case we can reinterpret the ring as giving a natural embedding of $C$ into a weighted projective space ${\mathbb{P}}$, such that $\omega_{C}={\mathcal{O}}_{\mathbb{P}}(2)|_{C}$ and ${\mathcal{L}}={\mathcal{O}}_{\mathbb{P}}(1)|_{C}/\text{torsion}$.
We hope these considerations to be of independent interest, a direct application to the moduli space of stable surfaces will be given in a subsequent paper.
@@ 2.A. Set-up
A curve $C$ is a Cohen-Macaulay projective scheme of pure dimension 1. It has a dualising sheaf $\omega_{C}$ and we denote by $p_{a}(C)$ the arithmetic genus of $C$, that is, $p_{a}(C)=1-\chi(\mathcal{O}_{C})$. The curve $C$ is Gorenstein if $\omega_{C}$ is invertible; if this is the case $K_{C}$ denotes a canonical divisor such that $\mathcal{O}_{C}(K_{C})\cong\omega_{C}$.
@@
A variety is a reduced (possibly reducible) Cohen–Macaulay scheme.
2.B. Rank one torsion-free sheaves on Gorenstein curves
A torsion-free sheaf on a curve $C$ is a coherent sheaf ${\mathcal{F}}$ such that $10pt({\mathcal{F}}_{x})=$ 1 for every closed point $x\in C$; equivalently ${\mathcal{F}}$ has no subsheaves supported on a finite set of points, i.e., ${\mathcal{F}}$ satisfies Serre condition $S_{1}$.
@@ 2.B. Rank one torsion-free sheaves on Gorenstein curves
[…]
@@
Let ${\mathcal{F}}$ be a rank one torsion free sheaf on a reduced Gorenstein curve. Then:
@@
    $\operatorname{Aut}({\mathcal{F}})=\operatorname{Aut}(\widetilde{\mathcal{F}})$. In particular, if $\widetilde{C}$ is connected then $\operatorname{Aut}({\mathcal{F}})={\mathbb{C}}^{*}$.
@@
Let $C$ be a reduced Gorenstein curve and let $A,B\subset C$ be subcurves without common components such that $C=A\cup B$. Then there is an exact sequence:
$0\to\omega_{B}\to\omega_{C}\to\omega_{C}{\raisebox{-0.1pt}{$|_{A}$}}\to 0$
@@
The proof follows by the arguments used in [CFHR99, Thm.1.1].
By Serre duality $H^{1}(C,{\mathcal{F}})^{\vee}\cong\operatorname{Hom}({\mathcal{F}},\omega_{C})$. Assume it is not zero and pick any nonzero map $\varphi\colon{\mathcal{F}}\to\omega_{C}$. By [CFHR99, Lemma 2.4], $\varphi$ comes from a generically surjective map ${\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}\to\omega_{B}$ for a subcurve $B\subseteq C$ and yields an injection ${\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}^{[{1}]}\hookrightarrow\omega_{B}$ whose cokernel has finite length. Therefore $\chi\left({\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}^{[{1}]}\right)\leq\chi(\omega_{B})$, and Remark 2.1 gives $\deg\left({\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}^{[{1}]}\right)\leq 2p_{a}(B)-2$, against the assumptions. $\square$
@@ 3. Reduced Gorenstein curves of genus two with ample canonical bundle
By Lemma 2.4 there is an exact sequence
(3.4) $0\to\omega_{C_{1}}\to\omega_{C}\to\omega_{C}|_{C_{2}}\to 0$
@@
$(iii)$ By Proposition 3.3 the curve $C$ is embedded in a weighted projective space $\mathbb{P}$ in such a way that the regular involution of $\mathbb{P}$ that sends $u$ to $-u$ and leaves the remaining variables invariant induces $\iota$ on $C$. $\square$
@@
Let $X$ be a variety (so in particular a Cohen–Macaulay scheme, cf. section 2.A) with an involution $\iota$ that acts non-trivially on every irreducible component of $X$ and let $\pi\colon X\to Y:=X/\iota$ be the quotient map. Then $\pi_{*}{\mathcal{O}}_{X}$ decomposes under the action of $\iota$ as a direct sum ${\mathcal{O}}_{Y}\oplus{\mathcal{F}}$, where ${\mathcal{F}}$ is a torsion-free sheaf of rank one, and $\iota$ acts on ${\mathcal{O}}_{Y}$ as the identity and on ${\mathcal{F}}$ as multiplication by $-1$. By relative duality we have $\pi_{*}\omega_{X}=\curly{H}om(\pi_{*}{\mathcal{O}}_{X},\omega_{Y})$ and the decomposition of $\pi_{*}{\mathcal{O}}_{X}$ induces a decomposition $\pi_{*}\omega_{X}=\omega_{Y}\oplus\curly{H}om({\mathcal{F}},\omega_{Y})$, where the first summand is $\iota$-invariant and the second one is anti-invariant. We call the corresponding linearisation of $\omega_{Y}$ “natural”. Note that when both $X$ and $Y$ are smooth the natural linearisation corresponds to the natural action on differential forms, for which invariant $n$-forms on $X$ correspond to pull-backs of $n$-forms on $Y$; note also that if the map $\pi$ is flat, namely ${\mathcal{F}}$ is a line bundle, then $\omega_{X}=\varphi^{*}(\omega_{Y}\otimes{\mathcal{F}}^{-1})$ and the opposite of the natural linearisation is the one induced by the fact that $\omega_{X}$ is a pull-back from $Y$.
@@
      $C$ is reducible iff $\sum_{i}(n_{i}+1)=6$ and all the $n_{i}$ are odd. The components of $C$ are smooth rational curves exchanged by $\iota$.
@@ 4. Generalised Gorenstein spin curves
Let $C$ be a reduced curve and $f\colon\widetilde{C}\to C$ a finite birational morphism. Then the normalisation of $C$ factors over $f$, which is therefore often called a partial normalisation. Note that a partial normalisation might not preserve the Gorenstein property in general: three general lines through the origin in ${\mathbb{C}}^{3}$ are not Gorenstein, but map birationally to three general lines through the origin in the plane, a Gorenstein singularity.
Example 4.2 —
$A_{2n}$:
$A_{2n-1}$:
where we consider ${\mathbb{C}}\llbracket x,x^{k}y\rrbracket$ as a subring of ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2})$, so that with $y_{k}=x^{k}y$ we get the desired relation $y_{k}^{2}-x^{k+2}=0$.
@@
Let $({\mathcal{L}},\mu)$ be a ggs structure on a reduced Gorenstein curve with only singularities of type $A_{m}$. Then there exists a (uniquely determined) partial normalisation $f\colon\widetilde{C}\to C$ and a regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on $\widetilde{C}$ such that
${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}\text{ and }{\mathcal{L}}^{[{2}]}=f_{*}\widetilde{\mathcal{L}}^{\operatorname{\otimes}2}.$
@@
It remains to show that (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ is a regular ggs structure on $\widetilde{C}$. We have seen in Example 4.2 that $\widetilde{C}$ is again Gorenstein, so we can apply relative duality for a finite morphism of Gorenstein schemes, e.g., [Sta23, Section 0AU3], which gives
$\operatorname{Hom}({\mathcal{L}}^{[2]},\omega_{C})=\operatorname{Hom}(f_{*}\widetilde{\mathcal{L}}^{\operatorname{\otimes}2},\omega_{C})\cong\operatorname{Hom}(\widetilde{\mathcal{L}}^{\operatorname{\otimes}2},\omega_{\widetilde{C}}),$
and clearly $0=\chi({\mathcal{L}})=\chi(\widetilde{\mathcal{L}})$. Restriction to the subset where $f$ is an isomorphism shows that $\varphi\colon{\mathcal{L}}^{[{2}]}\to\omega_{C}$ is generically an isomorphism if and only if the corresponding $\widetilde{\varphi}\colon\widetilde{\mathcal{L}}^{\operatorname{\otimes}2}\to\omega_{\widetilde{C}}$ is, completing the correspondence between ggs structures. $\square$
@@
Let us write ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$ for a regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on a partial normalisation $f\colon\widetilde{C}\to C$. Recall that by Corollary 3.9 the curve $\widetilde{C}$ is also Gorenstein. The composition $f\circ\varphi:\widetilde{C}\to Q$ is a degree 2 finite morphism, so it induces an involution $\widetilde{\iota}$ on $\widetilde{C}$ that lifts $\iota$.
We first observe that $\iota$ acts trivially on $\operatorname{Pic}^{0}(\widetilde{C})[2]$. Indeed, considering the induced action on $H^{1}({\mathcal{O}}_{\widetilde{C}})\cong H^{0}(\widetilde{C},\omega_{\widetilde{C}})$ and taking into account that $\widetilde{C}/\widetilde{\iota}$ has arithmetic genus 00 we see $\widetilde{\iota}$ acts on $H^{1}(\widetilde{C},{\mathcal{O}}_{\widetilde{C}})$ as multiplication by $-1$. Then the claim follows from the exponential sequence. It thus suffices to exhibit one regular $\widetilde{\iota}$-invariant ggs structure.
@@
Consider now $C$ of type A. By Corollary 3.9 if $\widetilde{C}$ is disconnected then it is the union of two smooth rational curves and we can take the unique ggs structure on each of them. If $\widetilde{C}$ is connected but reducible, then it is the union of two smooth rational curves, meeting either in two nodes or in an $A_{3}$ point, so it has genus 1 and we can take the trivial ggs structure ${\mathcal{O}}_{\widetilde{C}}$ on $\widetilde{C}$. If $\widetilde{C}\neq C$ is irreducible, then it has arithmetic genus at most one and admits either a trivial or a unique ggs structure, as above.
@@
If $g=2$ and $C$ is reducible, then by Proposition 4.5 and the geometry of the curves (Proposition 3.3) any ggs structure is irregular, so we consider the corresponding regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on a partial normalisation $\widetilde{C}$. If $\widetilde{C}$ is connected, then $h^{0}(\widetilde{\mathcal{L}})\leq 1$ by the above. If $\widetilde{C}$ is disconnected, then both components have genus at most one, so $h^{0}(C,{\mathcal{L}})\leq 2$ with equality if and only if $C=C_{1}\cup C_{2}$ is of type $B$ and ${\mathcal{L}}={\mathcal{O}}_{C_{1}}\oplus{\mathcal{O}}_{C_{2}}$ is the pushforward of the unique ggs structure with a section on the components, because the unique regular ggs structure on ${\mathbb{P}}^{1}$ does not have a section.
@@
The sheaf $f_{*}(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})$ is torsion-free and supported on $B$, so the last map in the above equation gives a surjective map ${\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]}\to f_{*}(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})$. Since both sheaves are torsion-free of rank one, the map is an isomorphism.
So we have $\chi({\mathcal{L}}{\raisebox{-0.1pt}{$|_{B}$}}^{[1]})=\chi({\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})=\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}+1-p_{a}(\widetilde{B}})$ and $2\deg(\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}})=\deg K_{\widetilde{C}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}$. Since there is an inclusion $\omega_{\widetilde{B}}\hookrightarrow\omega_{\widetilde{C}}$ (cf. Lemma 2.4), we get $\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}\geq p_{a}(\widetilde{B})-1\geq-1$, and if $\deg\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{\widetilde{B}}$}}=-1$ holds then $\widetilde{B}$ is smooth rational.

### v2

@@ 1. Introduction
Recall that a spin structure (or theta-characteristic) on a smooth curve $C$ is a line bundle $\mathcal{L}$} such that there is an isomorphism $\mu\colon{\mathcal{L}}\otimes{\mathcal{L}}\to\omega_{C}$. The interest in such structures originated in the classical study of theta functions in the 19th century and they were taken up again in 1970’s in the context of the study of moduli spaces of curves (see [Far12] for a recent survey). They also occur in the study of algebraic surfaces: if $X$ is a smooth projective surface and $C\in|K_{X}|$ is a smooth curve in the canonical linear system, then ${\mathcal{L}}=K_{X}|_{C}$ is a theta-characteristic by the adjunction formula and the section ring $R(C,{\mathcal{L}})$ is often a stepping stone to compute the canonical ring of $X$ by Reid’s hyperplane section principle.
In our study of the moduli space of stable surfaces, the need arose to consider a more general notion than what was available in the literature and which we hope to be of independent interest.
We define a generalised Gorenstein spin (ggs) structure on a Gorenstein curve $C$ as a torsion-free sheaf $\mathcal{L}$} of rank one at each generic point. such that $\chi({\mathcal{L}})=0$ together with a multiplication map $\mu\colon{\mathcal{L}}\operatorname{\otimes}{\mathcal{L}}\to\omega_{C}$ which is an isomorphism at the generic points.
Remark 1.1
Let us sketch our motivation for this definition. Let $X$ be a stable surface $K_{X}^{2}=1$ and $h^{0}(K_{X})=2$ such that $2K_{X}$ is Cartier. Then any canonical curve $C\in|K_{X}|$ is a Gorenstein curve of arithmetic genus two and with ample canonical bundle by adjuction. However, $\omega_{X}\operatorname{\otimes}{\mathcal{O}}_{C}$ might have torsion at the non-Gorenstein points of $X$. So we don’t get a theta characteristic in the classical sense, but dividing out the torsion results in a ggs structure on $C$.
The results of this paper are then an important puzzle piece in the complete classification of such surfaces $X$ in [CFPR24].
Inspired by Mumford’s work [Mum71], Harris [Ha82] considered the case of arbitrary curves with $\mathcal{L}$} a line bundle and the multiplication map being an isomorphism. The Gorenstein condition is then implicit and we call this a regular ggs structure. We will see in Section 4 that in favourable situations, the ggs structures on a curve $C$ are governed by regular ggs structures on partial normalisations.
In later studies of the (compactified) moduli space of spin curves, initiated by Cornalba [Cor89] and continued for example by Jarvis [Jar98] and Ludwig [Lud10], spin structures are only considered on (semi-)stable curves. In fact, on such curves their definition agrees with ours. [^1] The condition on the cokernel of the multiplication map in [Jar98, Definition 2.1.2] is automatic by Proposition 4.2, Example 2.1 and the degree condition.
@@
In our application to stable surfaces, Remark 1.1, we need to include both curves with cusps and curves with elliptic tails in our analysis. As explained in the comments after Definition 1.2 in [FS13], these two classes of curves cannot occur at the same time in a separated moduli problem for curves, so it is not clear how to deal with families in our setup. We have chosen not to pursue this question here.
Next, we apply these results to reduced Gorenstein curves of genus two with ample canonical bundle, the class which originally motivated our research. In Section 3 we give a complete description of such curves, showing that we can group them in two types, Type A and Type B, distinguished by the image of the bicanonical map (Proposition 3.3) and moreover we classify all possible ggs structures. Such curves are hyperelliptic in the sense of [BaBo24], i.e. they admit a $2:1$ morphism onto a rational (not necessarily irreducible) curve (see [Ho99] for the irreducible case).
@@
We then go on and analyse the half-canonical ring of ggs structures on such curves, that is,
@@
where $R_{2n}=H^{0}(C,nK_{C})$ and $R_{2n+1}=H^{0}(C,{\mathcal{L}}(nK_{C}))$. Note that this is not a standard section ring, because the multiplication of two elements of odd degree has to be defined via the multiplication map $\mu$.
Leaving the precise statements to Theorem 5.2 we prove the following:
Theorem
Let $({\mathcal{L}},\mu)$ be a ggs structure on a reduced Gorenstein curve of genus two with ample canonical bundle. Then $h^{0}({\mathcal{L}})=0,1,2$ and we have a complete description of the the half-canonical ring $R(C,\{{\mathcal{L}},\omega_{C}\})$ in terms of generators and relations according to the type of the curve and to the dimension $h^{0}({\mathcal{L}})$:
@@
In cases $A(0)$, $A(1)$, and $B(2)$ we get complete intersections, in cases $B(0)$ and $B(1)$ the rings admit a more complicated presentation in codimension six and four (see Table 2 for a complete description).
Note that in each case we can interpret the ring as giving a natural embedding of $C$ into a weighted projective space $\mathbb{P}$}, such that $\omega_{C}={\mathcal{O}}_{\mathbb{P}}(2)|_{C}$ and ${\mathcal{L}}={\mathcal{O}}_{\mathbb{P}}(1)|_{C}/\text{torsion}$.
@@ Acknowledgements
We thank the anonymous referee for some improvements of the presentation and interesting remarks.
@@ 2.A. Set-up
A variety is a reduced (possibly reducible) Cohen–Macaulay scheme.
2.A.1. Curves and partial normalisation
A curve $C$ is a Cohen–Macaulay projective scheme of pure dimension 1. It has a dualising sheaf $\omega_{C}$ and we denote by $p_{a}(C)$ the arithmetic genus of $C$, that is, $p_{a}(C)=1-\chi(\mathcal{O}_{C})$. The curve $C$ is Gorenstein if $\omega_{C}$ is invertible; if this is the case $K_{C}$ denotes a canonical divisor such that $\mathcal{O}_{C}(K_{C})\cong\omega_{C}$.
@@ 2.A.1. Curves and partial normalisation
Let $C$ be a reduced curve and $f\colon\widetilde{C}\to C$ a finite birational morphism. Then the normalisation of $C$ factors over $f$, which is therefore often called a partial normalisation.
Note that a partial normalisation might not preserve the Gorenstein property in general: three general lines through the origin in ${\mathbb{C}}^{3}$ are not Gorenstein, but map birationally to three general lines through the origin in the plane, a Gorenstein singularity.
For $A_{m}$ singularities these kinds of difficulties do not occur.
Example 2.1 —
$A_{2n}$:
$A_{2n-1}$:
where we consider ${\mathbb{C}}\llbracket x,x^{k}y\rrbracket$ as a subring of ${\mathbb{C}}\llbracket x,y\rrbracket/(y^{2}-x^{2})$, so that with $y_{k}=x^{k}y$ we get the desired relation $y_{k}^{2}-x^{2k+2}=0$.
2.A.2. Group action
Assume that we have a finite group $G$ acting on a projective variety $X$ and let $\mathcal{L}$} be a rank one torsion-free sheaf on $X$.
A linearisation of $G$ is an action of $G$ on $\mathcal{L}$} that lifts the action of $G$ on $X$. If $G$ is cyclic, generated by $g$, then the existence of a linearisation of $G$ is equivalent to say that there exists an isomorphism $g^{*}{\mathcal{L}}\cong{\mathcal{L}}$.
A linearisation of $\mathcal{L}$} induces a representation of $G$ on the vector space $H^{0}(X,{\mathcal{L}})$. Conversely, if $\mathcal{L}$} is ample, one can recover the action of $G$ by considering its action on $R(X,{\mathcal{L}})=\bigoplus_{d\in{\mathbb{N}}}H^{0}(X,{\mathcal{L}}^{\otimes d})$. We will apply these notions to projective varieties having an involution, which induces a decomposition of the space of sections in invariant and anti-invariants subspaces.
2.B. Weighted projective spaces
We recall some facts about weighted projective spaces, see [Dol, Fle00] for details.
Let $x_{0},\dots,x_{n}$ be coordinates on ${\mathbb{C}}^{n+1}$ and let $(w_{0},\dots,w_{n})$ be positive integers with $\gcd\{w_{1},\dots,\widehat{w}_{i},\dots,w_{n}\}=1$ for all $i=1,\dots,n$. Then the (well-formed) weighted projective space ${\mathbb{P}}(w_{0},\dots,w_{n})$ is defined as ${\mathbb{C}}^{n+1}/\sim$, where the equivalence relation $\sim$ is given, for $\lambda\in{\mathbb{C}}^{\ast}$, by
$(x_{0},\dots,x_{n})\sim(\lambda^{w_{0}}x_{0},\dots,\lambda^{w_{n}}x_{n}).$
[…]
$0\to{\mathcal{F}}_{c}\to\dots\to{\mathcal{F}}_{0}\to{\mathcal{O}}_{X}\to 0$
[…]
2.C. Rank one torsion-free sheaves on Gorenstein curves
A torsion-free sheaf on a curve $C$ is a coherent sheaf $\mathcal{F}$} such that $0pt({\mathcal{F}}_{x})=1$ for every closed point $x\in C$; equivalently $\mathcal{F}$} has no subsheaves supported on a finite set of points, i.e., $\mathcal{F}$} satisfies Serre condition $S_{1}$.
@@ 2.C. Rank one torsion-free sheaves on Gorenstein curves
${\mathcal{F}}^{[m]}=\left({\mathcal{F}}^{\operatorname{\otimes}m}\right)^{\vee\vee}=\left({\mathcal{F}}^{\operatorname{\otimes}m}\right)/\langle\text{0-dimensional torsion}\rangle.$
@@
Let $\mathcal{F}$} be a rank one torsion free sheaf on a reduced Gorenstein curve and assume that all the partial normalisations of $C$ are Gorenstein. Then:
@@
[…]
@@
Let $C$ be a reduced Gorenstein curve and let $A,B\subset C$ be subcurves without common components such that $C=A\cup B$. Then there are a natural isomorphism ${\mathcal{I}}_{A}\omega_{C}\cong\omega_{B}$ and an exact sequence:
$0\to\omega_{B}\to\omega_{C}\to\omega_{C}{\raisebox{-0.1pt}{$|_{A}$}}\to 0.$
@@
The proof follows by the arguments used in [CFHR99, Thm.1.1]. By Serre duality $H^{1}(C,{\mathcal{F}})^{\vee}\cong\operatorname{Hom}({\mathcal{F}},\omega_{C})$. Assume it is not zero and pick any nonzero map $\varphi\colon{\mathcal{F}}\to\omega_{C}$. By [CFHR99, Lemma 2.4], $\varphi$ comes from a generically surjective map ${\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}\to\omega_{B}$ for a subcurve $B\subseteq C$ and yields an injection ${\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}^{[{1}]}\hookrightarrow\omega_{B}$ whose cokernel has finite length. Therefore $\chi\left({\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}^{[{1}]}\right)\leq\chi(\omega_{B})$, and Remark 2.2 gives $\deg\left({\mathcal{F}}{\raisebox{-0.1pt}{$|_{B}$}}^{[{1}]}\right)\leq 2p_{a}(B)-2$, against the assumptions. $\square$
@@ 3. Reduced Gorenstein curves of genus two with ample canonical bundle
By Lemma 2.5 we have an isomorphism ${\mathcal{I}}_{C_{2}}\omega_{C}\cong\omega_{C_{1}}$ and there is an exact sequence
(3.4) $0\to\omega_{C_{1}}\to\omega_{C}\to\omega_{C}|_{C_{2}}\to 0.$
@@
$(iii)$ By Proposition 3.3 the curve $C$ is embedded in a weighted projective space $\mathbb{P}$ in such a way that the regular involution of $\mathbb{P}$ that sends $u$ to u-u and leaves the remaining variables invariant induces $\iota$ on $C$. $\square$
@@
Let $X$ be a projective variety, in particular reduced and Cohen–Macaulay by our conventions, with an involution $\iota$ that acts non-trivially on every irreducible component of $X$ and let $\pi\colon X\to Y:=X/\iota$ be the quotient map. Then $\pi_{*}{\mathcal{O}}_{X}$ decomposes under the action of $\iota$ as a direct sum ${\mathcal{O}}_{Y}\oplus{\mathcal{F}}$, where $\mathcal{F}$} is a torsion-free sheaf of rank one, and $\iota$ acts on ${\mathcal{O}}_{Y}$ as the identity and on $\mathcal{F}$} as multiplication by 1-1. By relative duality we have $\pi_{*}\omega_{X}=\curly{H}om(\pi_{*}{\mathcal{O}}_{X},\omega_{Y})$ and the decomposition of $\pi_{*}{\mathcal{O}}_{X}$ induces a decomposition $\pi_{*}\omega_{X}=\omega_{Y}\oplus\curly{H}om({\mathcal{F}},\omega_{Y})$, where the first summand is $\iota$-invariant and the second one is anti-invariant. We call the corresponding linearisation of $\omega_{X}$ natural. Note that when both $X$ and $Y$ are smooth the natural linearisation corresponds to the natural action on differential forms, for which invariant $n$-forms on $X$ correspond to pull-backs of $n$-forms on $Y$; note also that if the map $\pi$ is flat, namely $\mathcal{F}$} is a line bundle, then $\omega_{X}=\varphi^{*}(\omega_{Y}\otimes{\mathcal{F}}^{-1})$ and the opposite of the natural linearisation is the one induced by the fact that $\omega_{X}$ is a pull-back from $Y$.
@@
      $C$ is reducible iff $\sum_{i}(n_{i}+1)=6$ and all the $n_{i}$ are odd. In this case, the components of $C$ are smooth rational curves exchanged by $\iota$.
@@ 4. Generalised Gorenstein spin curves
[…]
@@
[…]
@@
[…]
@@
Consider now $C$ of type A. By Corollary 3.9 if $\widetilde{C}$ is disconnected then it is the union of two smooth rational curves and we can take the unique ggs structure on each of them. If $\widetilde{C}$ is connected but reducible, then it is the union of two smooth rational curves, meeting either in two nodes or in an $A_{3}$ point, so it has genus 1 and we can take the trivial ggs structure ${\mathcal{O}}_{\widetilde{C}}$ on $\widetilde{C}$. If $\widetilde{C}\neq C$ is irreducible, then it has arithmetic genus at most one and admits either a trivial or a unique regular ggs structure, as above.
@@
[…]
@@
[…]